OpenStudy (garretttyre05):

Which shows the expressions in the order they would appear on a number line from least to greatest?

OpenStudy (taylor0402):

Where is it

OpenStudy (garretttyre05):

Question 1 (Multiple Choice Worth 5 points) (01.07A MC) Which shows the expressions in the order they would appear on a number line from least to greatest? 2 to the power of 3, square root of 5, square root of 20, square root of 11, 11 over 9 2 to the power of 3, square root of 11, 11 over 9, square root of 20, square root of 11 11 over 9, square root of 5, square root of 11, square root of 20, 2 to the power of 3 11 over 9, 2 to the power of 3, square root of 5, square root of 20, square root of 11

OpenStudy (garretttyre05):

Question 1 (Multiple Choice Worth 5 points) (01.07A MC) Which shows the expressions in the order they would appear on a number line from least to greatest? A. 2 to the power of 3, square root of 5, square root of 20, square root of 11, 11 over 9 B. 2 to the power of 3, square root of 11, 11 over 9, square root of 20, square root of 11 C. 11 over 9, square root of 5, square root of 11, square root of 20, 2 to the power of 3 D. 11 over 9, 2 to the power of 3, square root of 5, square root of 20, square root of 11

OpenStudy (anonymous):

Convert them all to decimal, and check.

OpenStudy (garretttyre05):

do u have any idea?

OpenStudy (texaschic101):

2^3 = 8................sq rt 11 = 3.32 sq rt 5 = 2.24........11/9 = 1.22 sq rt 20 = 4.47 can u take it from here ?

OpenStudy (garretttyre05):

so it would be c?

OpenStudy (texaschic101):

yes

OpenStudy (garretttyre05):

ok thx